• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 20:27:56 on July 21, 2006

    Entry number 174761

    This entry is a followup to: 174717

    keyword=BCMs: new amplifier on downstream cavity seems not to be amplifiying

    It appears that the signals from our BCMs (now equipped with the new amplifiers courtesy of John Musson and Emily McCullough) show the following: the upstream cavity appears amplified appropriately (calibrations to OLO2 coming later in the plan), but the downstream cavity appears unamplified (i.e. looks to be giving a response similar, maybe the same, as without the amplifier). Note that both cavities present identical responses (voltages upstairs, after the "10 kHz RMS-DC Converter" box) when there is no beam - but the difference in amplitude of response is readily apparent with even a little beam going through.

    Emily suspects it may be due to a broken filter in the amplifier box (in the hall), but would need time with John Musson to trouble-shoot.

    We can live with this anomaly over the weekend for the CSR tests, but should definitely plan to fix this during the changeover from CSR to E05-103 on Monday.

    A copy of this log entry has been emailed to: musson@jlab.org,emily@Jlab.org